OpenStudy (btaylor):

Multivariable Calculus: Find the volume of an ellipsoid with semiaxes a, b, and c by double integration, using either rectangular or polar coordinates. I have the equation for the general (rectangular) form but would rather use polar to avoid crazy trig substitution. Thanks in advance!! @thomaster @iambatman @e.mccormick @campbell_st

OpenStudy (btaylor):

Right now I have the double integral in rectangular form as: \[\int\limits_{-a}^{a} \left(\int\limits_{-b\sqrt{1-x^2/a^2}}^{b\sqrt{1-x^2/a^2}} c\sqrt{1-\frac{x^2}{a^2}-\frac{y^2}{b^2}}dy\right)dx\] Parentheses are just to make the limits of integration clear

OpenStudy (perl):

ok your ellipsoid is (x/a)^2 + (y/b)^2 + (z/c)^2 = 1 and the volume should be V = 4/3 Pi * a*b*c

OpenStudy (btaylor):

Yes, but I need to derive that using the double integral.

OpenStudy (btaylor):

I wish I could just write down the answer and get credit

OpenStudy (perl):

but we need an expression in terms of x^2 + y^2

OpenStudy (perl):

I don't see an easy way to do it using polar coordinates. but rectangular should work , or you might be able to use spherical coordinates

OpenStudy (perl):

i agree with your integral, you just need to multiply it by 2 since you are finding the volume of the top half of the ellipsoid

OpenStudy (perl):

would you like me to integrate that?

OpenStudy (btaylor):

Yes, that would be awesome. I think I figured out how to set it up in polar form, but am struggling to integrate \[\frac{1}{\frac{\cos^2 \theta}{a^2} + \frac{\sin^2 \theta}{b^2}}\]

OpenStudy (perl):

the double rectangular integral works*

OpenStudy (perl):

make sure to put 2 * in front

OpenStudy (btaylor):

Can you help me get started on the double rectangular integral?

OpenStudy (perl):

I used a computer algebra system to check

OpenStudy (perl):

its pretty hairy to integrate that

OpenStudy (btaylor):

I know. but i have to. Which is why I'm trying to do it in polar.

OpenStudy (perl):

right, the problem with the polar is that a^2 and b^2

OpenStudy (btaylor):

I feel like if I can integrate that last part it will work. I'm gonna try if wolfram can do a step-by-step

OpenStudy (btaylor):

not working...

OpenStudy (btaylor):

Thanks. I have to go now, but will check back in to see what you came up with

OpenStudy (perl):

I can solve this using triple coordinates and Jacobian transformation. First use the transformation x= a*u , y = b*v , z = c*w then J = Determinant{ {dx/du, 0, 0 }, {0, dy/dv, 0}, {0, 0, dz/dw} } = DET { {a, 0, 0 }, {0, b, 0}, {0, 0, c} } = abc Therefore V = int(int(int dx) dy)dz over the ellipsoid (x^2/a^2 + y^2/b^2 + z^2/c^2 <= 1) = (abc) int(int(int du)dv)dw over the sphere (u^2 + v^2 + w^2 <= 1) But volume of a sphere = 4/3 Pi r^3 = 4Pi/3 r^3 (abc) here r = 1 = (4Pi / 3) abc

OpenStudy (anonymous):

that looks correct to me